OpenStudy (anonymous):

SOLVE FOR X (1/3X)=6

OpenStudy (anonymous):

SOLVE FOR X (1/3 X)=6

OpenStudy (anonymous):

the solution to \[\frac{1}{3}x=6\]is found by multiplying both side by 3 to clear the denominator\[3*\frac{1}{3}x=3*6\]so\[x=18\]